# Reasoning in the Many: Dynamics of Collaboration Across Domains and Scales Towards Multi-Agent Reasoning Systems for Collaborative Expertise Delegation: An Exploratory Design Study

### v1

@@ 1 Introduction
The recent evolution of large language and reasoning models (LLMs/LRMs) Yang et al. (2025); Jaech et al. (2024) has spurred parallel investigations into machine collective intelligence. Contemporary research has developed artificial analogs of human collaboration patterns through techniques such as multi-agent debate frameworks and workflow orchestration Li et al. (2024c); Du et al. (2024); Liang et al. (2024). These approaches primarily focus on either enhancing individual model performance through collective verification processes or establishing general-purpose problem-solving workflow pipelines.
@@
Another dimension concerning the multi-agent system design is the collaboration paradigm—namely, the mechanism governing how specialist agents interact in a multi-agent system. Currently, the collaboration paradigm predominantly used in recent studies could be categorized into two kinds: (1) Diversity-Driven Perspective Integration, where agents, often embodying different viewpoints or roles, are encouraged to generate diverse responses to enrich the solution space Wang et al. (2024b); Chen et al. (2024b). (2) Structured Workflow Cooperation, where different agents are assigned distinct sub-tasks within a predefined pipeline to collaboratively construct a solution Chen et al. (2024c); Hong et al. (2024). To understand the preference of collaboration paradigm in collaborative expertise specialization, we design comparative experiments which unveil the performance differences between paradigms. Our observations reveal a consistent advantage for diversity-driven collaboration over structured workflow collaboration, suggesting the superiority of the diversity-driven paradigm. Detailed analysis regarding viewpoint diversity are also conducted to study the impact of agent response diversity in multi-agent system, where a higher diversity could indicate a better performance.
@@ 3.1 Problem Setup
where stands for the final answer generated by the system. represents the answer of agent i, stands for the communication protocol manually customized by the design of the system which aggregate the answer of each agents into the final answer. Typically, it could be majority vote, debate, etc Kaesberg et al. (2025); Liu et al. (2024a). In our specific setup, we adopt a sequential processing communication mechanism inspired by Qian et al. (2024) to prevent context explosion Liu et al. (2024b); Xu et al. (2024c). In this mechanism, for , agent receives the complete output generated by the immediately preceding agent . In contrast, from the preceding agents {}, receives only the final answers. The detailed communication algorithm could be found in Algorithm 1.
@@ 3.4 General Experiment Setup
As detailed in Section 3.2, we select 4 representative domains from MMLU-pro to investigate the effects of collaborative expertise specialization. To be consistent with all experiments, we utilize DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI et al. (2025) as the foundational model for all experiments. Each agent is initialized with its specific expert description and responsibilities via its system prompt, while the task instance is provided through the user prompt. The detailed prompts could be found in Appendix D. All experiments adopt accuracy as the evaluation metric.
@@ 4.1 Setup
Considering the primary principle of multi-agent reasoning system is to incorporate more diverse agent viewpoints and integrate them in the final answer Liang et al. (2024), in our experiments, we adopt diversity-driven collaboration paradigm where we distribute each agent with a specific domain expert configuration and instruct them to generate responses based on their expertise. At this stage, we fix the size of the multi-agent reasoning system to be 3 for controllable computational cost.
@@ 5.1 Setup
To ensure a plausible, accurate generation of expert role descriptions, we employ GPT-4o OpenAI et al. (2024). The model was prompted with the target expert domain, the group size, and the specific collaboration paradigm as inputs. The detailed prompts utilized for this automated role generation process are provided in Appendix B.1.
@@ 5.3 Analysis on Response Diversity
To quantitatively characterize how the collaboration paradigm influences the diversity of agent contributions, we further design a response diversity analysis. We leverage semantic embeddings derived from Sentence-BERT Reimers and Gurevych (2019). For each task instance solved by the multi-agent system, we generate embeddings for the textual output of each agent in the sequence. We then measure the internal diversity of the system’s responses by calculating the pairwise cosine similarity between the embeddings of outputs from different agents. This provides a measure of how semantically distinct the contributions are at different stages.
The distributions of these pairwise similarity scores for both the diversity-driven and structured workflow paradigms are presented in Figure 5. The results clearly indicate that, the pairwise cosine similarity values are consistently lower for the diversity-driven collaboration paradigm compared to the structured workflow paradigm. This finding demonstrates that the diversity-driven approach, which emphasizes fine-grained expertise, fosters greater semantic diversity among agent responses throughout the collaborative reasoning process, further confirming that enhancing the diversity of perspectives within a multi-agent system would be a key factor in improving its overall reasoning performance.
@@ 6.1 Setup
Generating coherent and appropriately specialized expert role configurations for these larger systems requires extending the initial configurations of the 3 agent system and we continue to leverage GPT-4o for this purpose. The detailed prompts employed for this role augmentation process are provided in Appendix B.2
@@ 7 Conclusions
In conclusion, this paper systematically investigate the impact of multi-agent system expertise specialization on collective reasoning intelligence across three key dimensions: expertise-domain alignment, collaboration paradigm, and system scale. Our empirical results demonstrate that the effectiveness of aligning agent expertise with task domains is highly contingent on the reasoning type required. Furthermore, we find that collaboration paradigms emphasizing the integration of diverse perspectives significantly outperform structured task decomposition, with response diversity playing a crucial role. Finally, while increasing the number of expert agents generally boosts reasoning performance, our analysis indicates diminishing returns concerning the performance gain relative to computational cost. These findings provide actionable insights for designing specialized multi-agent reasoning systems in future researches and underscore the critical need for developing more efficient coordination protocol as systems scale.
@@ Appendix A Agent Communication Algorithm
Appendix B Expert Generation Prompts
B.1 Primary Expert Generation Prompts
B.2 Expert Augmentation Process
Appendix C Social Group Role Examples
Appendix D Role System Prompt

### v2

# Towards Multi-Agent Reasoning Systems for Collaborative Expertise Delegation: An Exploratory Design Study

@@ 1 Introduction
The recent evolution of large language and reasoning models (LLMs/LRMs) Yang et al. (2025); Jaech et al. (2024); Team et al. (2025) has spurred parallel investigations into machine collective intelligence. Contemporary research has developed artificial analogs of human collaboration patterns through techniques such as multi-agent debate frameworks and workflow orchestration Li et al. (2024c); Du et al. (2024); Liang et al. (2024). These approaches primarily focus on either enhancing individual model performance through collective verification processes or establishing general-purpose problem-solving workflow pipelines.
@@
Another dimension concerning the multi-agent system design is the collaboration paradigm—namely, the mechanism governing how specialist agents interact in a multi-agent system. Currently, the collaboration paradigm predominantly used in recent studies could be categorized into two kinds: (1) Diversity-Driven Perspective Integration, where agents, often embodying different viewpoints or roles, are encouraged to generate diverse responses to enrich the solution space Wang et al. (2024b); Chen et al. (2024b); Hu et al. (2025). (2) Structured Workflow Cooperation, where different agents are assigned distinct sub-tasks within a predefined pipeline to collaboratively construct a solution Chen et al. (2024c); Hong et al. (2024). To understand the preference of collaboration paradigm in collaborative expertise specialization, we design comparative experiments which unveil the performance differences between paradigms. Our observations reveal a consistent advantage for diversity-driven collaboration over structured workflow collaboration, suggesting the superiority of the diversity-driven paradigm. Detailed analysis regarding viewpoint diversity are also conducted to study the impact of agent response diversity in multi-agent system, where a higher diversity could indicate a better performance.
@@ 3.1 Problem Setup
where stands for the final answer generated by the system. represents the answer of agent i, stands for the communication protocol manually customized by the design of the system which aggregate the answer of each agents into the final answer. Typically, it could be majority vote, debate, etc Kaesberg et al. (2025); Liu et al. (2024a). In our specific setup, we adopt a sequential processing communication mechanism inspired by Qian et al. (2024) to prevent context explosion Liu et al. (2024b); Xu et al. (2024c). In this mechanism, for , agent receives the complete output generated by the immediately preceding agent . In contrast, from the preceding agents {}, receives only the final answers. The detailed communication algorithm could be found in Appendix A algorithm 1.
@@ 3.4 General Experiment Setup
As detailed in Section 3.2, we select 4 representative domains from MMLU-pro to investigate the effects of collaborative expertise specialization. To be consistent with all experiments, we utilize DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI et al. (2025) as the foundational model for all experiments. Each agent is initialized with its specific expert description and responsibilities via its system prompt, while the task instance is provided through the user prompt. The detailed prompts could be found in Appendix B. All experiments adopt accuracy as the evaluation metric.
@@ 4.1 Setup
Considering the primary principle of multi-agent reasoning system is to incorporate more diverse agent viewpoints and integrate them in the final answer Liang et al. (2024), in our experiments, we adopt diversity-driven collaboration paradigm where we distribute each agent with a specific domain expert configuration and instruct them to generate responses based on their expertise. At this stage, we fix the size of the multi-agent reasoning system to be 3 for controllable computational cost. We employ GPT-4o OpenAI et al. (2024). The model is prompted with the target expert domain, the group size as inputs. The detailed prompts utilized for this automated role generation process are provided in Appendix C.
@@ 5.1 Setup
To ensure a plausible, accurate generation of expert role descriptions, we continue to employ GPT-4o with collaboration paradigm as extra input.
@@ 5.3 Analysis on Response Diversity
To quantitatively characterize how the collaboration paradigm influences the diversity of agent contributions, we further design a response diversity analysis. We leverage semantic embeddings derived from Sentence-BERT Reimers and Gurevych (2019). For each task instance solved by the multi-agent system, we generate embeddings for the output of each agent. We then measure the internal diversity of the system’s responses by calculating the pairwise cosine similarity between the embeddings of outputs from different agents. This provides a measure of how semantically distinct the contributions are at different stages.
The distributions of these pairwise similarity scores for both the diversity-driven and structured workflow paradigms are presented in Figure 5. The results clearly indicate that, the pairwise cosine similarity values are consistently lower for the diversity-driven collaboration paradigm compared to the structured workflow paradigm. This finding demonstrates that the diversity-driven approach, which emphasizes fine-grained expertise, fosters greater semantic diversity among agent responses throughout the collaborative reasoning, further confirming that enhancing the diversity of perspectives within a multi-agent system would be a key factor in improving its overall reasoning performance.
@@ 6.1 Setup
Generating coherent and appropriately specialized expert role configurations for these larger systems requires extending the initial configurations of the 3 agent system and we continue to leverage GPT-4o for this purpose. The detailed prompts employed for this role augmentation process are provided in Appendix C
@@ 7 Conclusions
In conclusion, this paper systematically investigate the impact of multi-agent system expertise specialization on collective reasoning intelligence across three key dimensions: expertise-domain alignment, collaboration paradigm, and system scale. Our empirical results demonstrate that the effectiveness of aligning agent expertise with task domains is highly contingent on the reasoning type required. Furthermore, we find that collaboration paradigms emphasizing the integration of diverse perspectives outperform structured task decomposition, with response diversity playing a crucial role. Finally, a scaling law behind collaborative expertise specialization in multi-agent system is observed where stacking more experts could result in better reasoning performance. These findings provide actionable insights for designing specialized multi-agent reasoning systems in future researches and underscore the critical need for developing more efficient coordination protocol as systems scale.
@@ Appendix A Agent Communication Algorithm
Appendix B Role System Prompt
Appendix C Expert Generation Prompts
C.1 Primary Expert Generation Prompts
C.2 Expert Augmentation Process
Appendix D Social Group Role Examples